OpenStudy (anonymous):

Note: You may use the following formulas to help you complete the data table: °C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273 Planet Name Average Temperature (°F) Average Temperature (°C) Average Temperature (°K) Mercury Venus Earth Mars Jupiter Saturn Uranus Neptune

OpenStudy (anonymous):

please someone help me! I will fan you even give yo a medal!

OpenStudy (anonymous):

@ganeshie8 @kausarsalley @k.rajabhishek help me fill the chart please i dont understand how to use this C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273